# Enhancing Multimodal Large Language Models with Vision Detection Models: An Empirical Study

### v2

@@ E.1 Modification on the GQA Benchmark
Thus, we make modifications to the GQA benchmark. We select only a subset of the evaluation dataset, including samples that only require yes or no answers, as well as those involving choices (questions containing “or”). For these samples, their correct answers can be chosen from a limited set of options, eliminating the possibility of evaluated models providing correct but non-matching answers, which leads to more accurate evaluation outcomes. After filtering, the remaining number of samples is 5,677, approximately half of the original evaluation dataset. We name the modified evaluation benchmark as GQA*.
@@ E.2 MME Benchmark in Table 4
Table 16: Comparison of SOTA MLLMs on the MME benchmark.
@@
BLIP-2 1293.8 290.0
InstructBLIP 1212.8 291.8
Qwen-vl 1400.6 266.4
Qwen-VL-Chat 1487.5 360.7
SPHINX 1476.1 322.2
@@
In Table 16, we present benchmark scores for RBI models and FTBI models on MME-Perception and MME-Cognition. The value reveals a significant enhancement in scores for both models specifically on MME-Cognition. This notable enhancement can be ascribed to the inclusion of supplementary OCR information, addressing a multitude of questions within MME-Cognition that pertain to textual content embedded within images.
@@
Furthermore, concerning the MME-Perception benchmark, our models exhibit some fluctuations in scores. Nonetheless, it is noteworthy that the scores for FTBI models surpass those for RBI models, which underscores that our third fusion approach better preserves the original capabilities of MLLM.
E.3 Performance on the VALSE Benchmark
@@
VALSE[41](Vision And Language Structured Evaluation) is a zero-shot benchmark designed to test the visual-linguistic grounding capabilities of general-purpose pre-trained vision-language models on specific linguistic phenomena. It can assess many capabilities of MLLMs, including six aspects: existence, plurality, counting, spatial relations, actions, and entity co-reference.
@@
In VALSE, a valid instance consists of an image, a caption, and a modified caption called a ‘foil’ that exemplifies a specific linguistic phenomenon. The tested model is required to distinguish between real captions and foils. VALSE employs four metrics to evaluate the model’s performance: overall accuracy ($acc$) on all classes (foil and correct); precision ($p_{c}$) measuring how well models identify the correct examples; foil precision($p_{f}$) measuring how well foiled cases are identified; and pairwise ranking accuracy ($acc_{r}$), which measures whether the image-sentence alignment score is greater for a correct image-text pair than for its foiled pair. $acc_{r}$ is more permissive than $acc$ as it accepts model predictions if the score for a foil is lower than the caption’s score.
@@
Due to the inability of LLaVA-1.5 and our models to directly output “cross_relationship_score” as the image-sentence alignment score like models such as LXMERT, we modify the computation of $acc_{r}$, $acc$, $p_{c}$ and $p_{f}$ following the approach outlined in “lxmert_valse_eval.py” (https://github.com/Heidelberg-NLP/VALSE/blob/main/lxmert_valse_eval.py) as follows:
@@
Table 17: Comparison between FTBI-7B and LLaVA-1.5-7B on the VALSE benchmark.
Existence Plurality Counting_hard Counting_small
$acc_{r}$ LLaVA-v1.5-7B 69.9 13.4 35.9 35.6
FTBI-7b 70.5 17.6 46.1 51.6
$acc$ LLaVA-v1.5-7B 84.0 56.2 64.6 66.9
FTBI-7b 84.1 58.1 71.1 74.4
$min(p_{c},p_{f})$ LLaVA-v1.5-7B 73.7 16.0 52.1 45.7
FTBI-7b 71.5 22.1 69.5 65.3
Counting_adversarial Relations Action Replacement Actant Swap
$acc_{r}$ LLaVA-v1.5-7B 25.2 4.7 34.3 10.3
FTBI-7b 36.3 8.2 37.4 19.2
$acc$ LLaVA-v1.5-7B 55.6 52.0 66.4 53.1
FTBI-7b 64.8 53.4 67.6 57.3
$min(p_{c},p_{f})$ LLaVA-v1.5-7B 39.8 7.5 43.8 16.7
FTBI-7b 59.5 14.6 52.9 30.2
Coreference Coreference_hard Foil_it
$acc_{r}$ LLaVA-v1.5-7B 5.2 4.8 50.5
FTBI-7b 20.2 18.3 63.0
$acc$ LLaVA-v1.5-7B 52.3 52.4 75.1
FTBI-7b 58.6 55.8 81.4
$min(p_{c},p_{f})$ LLaVA-v1.5-7B 6.4 4.8 53.5
FTBI-7b 24.0 20.2 66.6
The results are presented in Table 17. It can be observed that FTBI-7B outperforms LLaVA-v1.5-7B in all aspects, which indicates that our model is more sensitive to foiled instances and has better capabilities in visual grounding. It also emphasizes that the incorporation of the object detection module has indeed enhanced the object detection proficiency of the MLLM.
E.4 Applying Weights on the Aggregated Score
Table 18: The aggregated score with different weights.
$Weight=0.5$ $Weight=0$
InstructBLIP 0.285 0.345
Qwen-VL 0.574 0.668
Qwen-VL-Chat 0.765 0.778
SPHINX 0.713 0.665
LLaVA-1.5-7B 0.769 0.771
RBI-7B (ours) 0.830 0.832
FTBI-7B (ours) 0.870 0.882
LLaVA-1.5-13B 0.839 0.834
RBI-13B (ours) 0.896 0.898
FTBI-13B (ours) 0.946 0.954
The purpose of introducing the aggregated score is to provide a rough comparison of our models’ overall performance with other MLLMs, especially LLaVA-1.5. In our analysis, we don’t place too much emphasis on the comparison of this aggregated score. Instead, we focus more on comparing individual benchmarks and analyzing the performance changes of our models.
It’s possible for the aggregated score to introduce unfairness due to the varying significance of different benchmarks. To enhance fairness, we introduce weights to calculate the aggregated score based on the importance of each benchmark. Specifically, we assign weights of 0.5 or 0 to less significant benchmarks MM-Bench and Seed-Bench, while setting the weights of other benchmarks to 1.
According to Table 18, with the introduction of low weights for less significant benchmarks, our models still exhibit higher scores, demonstrating superior overall performance.

### v3

@@ E.1 Modification on the GQA Benchmark
Thus, we make modifications to the GQA benchmark. We select only a subset of the evaluation dataset, including samples that only require yes or no answers, as well as those involving choices (questions containing “or”). For these samples, the answer can be chosen from a limited set of options, eliminating the possibility of models providing correct but non-matching answers, which leads to more accurate evaluation outcomes. After filtering, the remaining number of samples is 5,677, approximately half of the original evaluation dataset. We name the modified evaluation benchmark as GQA*.
@@ E.2 MME Benchmark in Table 1
Table 19: Performance of TFI models, RBI models, and FTBI models on the MME benchmark.
@@
In Table 19, we present benchmark scores for TFI models, RBI models, and FTBI models on MME-Perception and MME-Cognition. According to the table, it reveals a significant enhancement in scores for both models on MME-Cognition. This notable enhancement can be ascribed to the infusion of supplementary OCR information, addressing a multitude of questions within MME-Cognition that pertain to textual content embedded within images.
@@
Furthermore, concerning the MME-Perception benchmark, our models exhibit some fluctuations in scores. Nonetheless, it is noteworthy that the scores for FTBI models surpass those for TFI models and RBI models, which underscores that the fine-tuning approach better preserves the original image understanding capabilities of MLLMs.
E.3 Performance on DocumentVQA Benchmarks
In this subsection, we evaluate our models on two well-known DocumentVQA benchmarks, DocVQA (Mathew et al., 2021) and InfographicVQA(Mathew et al., 2022). These benchmarks are specifically designed for visual question answering tasks where questions are answered using text within the document images. Their datasets provide OCR transcriptions and ground truth answers, enabling the evaluation of models in interpreting and extracting information from documents.
The results are presented in the two tables below. The first table compares the performance of the TFI, RBI, and FTBI models on the DocVQA and InfographicVQA benchmarks. The second table compares the performance of the RBI and FTBI models on the same benchmarks without incorporating detection information during inference.
Table 20: Performance of the TFI, RBI, and FTBI models on DocVQA and InfographicVQA.
Model DocVQA InfographicVQA
LLaVA-1.5-7B 19.4 18.8
LLaVA-1.5-7B-TFI 35.3 21.0
LLaVA-1.5-7B-RBI 35.7 20.9
LLaVA-1.5-7B-FTBI 35.9 21.3
LLaVA-1.5-13B 20.6 20.7
LLaVA-1.5-13B-TFI 35.5 22.1
LLaVA-1.5-13B-RBI 37.9 23.3
LLaVA-1.5-13B-FTBI 38.5 24.2
Table 21: Performance of the TFI, RBI, and FTBI models on DocVQA and InfographicVQA (without detection information being input during inference).
Model DocVQA InfographicVQA
LLaVA-1.5-7B 19.4 18.8
LLaVA-1.5-7B-RBI w/o DI 17.3 17.8
LLaVA-1.5-7B-FTBI w/o DI 19.4 18.7
LLaVA-1.5-13B 20.6 20.7
LLaVA-1.5-13B-RBI w/o DI 18.6 20.1
LLaVA-1.5-13B-FTBI w/o DI 20.6 20.9
As shown in Table 20, the deployment of detection models, particularly the OCR model, leads to a significant score improvement on DocVQA. Furthermore, models with adaptive training noticeably outperform training-free models . Specifically, the FTBI models surpass the RBI models, which in turn outperforms the TFI models. This suggests that the adaptive training enables MLLMs to better leverage the input detection information, resulting in improved performance.
Table 21 presents a comparison between the RBI models and the FTBI models in the absence of infused detection information. As shown, the performance of the RBI models is significantly inferior to that of the FTBI models. While the FTBI models, without detection information, perform similarly to the original LLaVA-1.5. This demonstrates that the fine-tuning strategy allows MLLMs to effectively balance the weights between the image encoder output and textual detection information, thereby preserving the comprehensive VQA capabilities. These results are consistent with the findings on the main page.
E.4 Performance on the VALSE Benchmark
@@
VALSE (Parcalabescu et al., 2022) (Vision And Language Structured Evaluation) is a zero-shot benchmark designed to test the visual-linguistic grounding capabilities of general-purpose vision-language models on specific linguistic phenomena. It assesses many capabilities of MLLMs, including six aspects: existence, plurality, counting, spatial relations, actions, and entity co-reference. In this subsection, we will evaluate the performance of LLaVA-1.5, TFI-7B, and FTBI-7B on the VALSE benchmark and compare their results. This analysis will further validate our conclusion on the main page: the fine-tuning strategy enables MLLMs to better understand the input texual detection information compared to the training-free approach.
@@
In VALSE, a valid instance consists of an image, a caption, and a modified caption called a ‘foil’ that exemplifies a specific linguistic phenomenon. The tested model is required to distinguish between real captions and foils. VALSE employs four metrics to evaluate the model’s performance: overall accuracy ($acc$) on all classes (foil and correct); precision ($p_{c}$) measuring how well models identify the correct examples; foil precision($p_{f}$) measuring how well foiled cases are identified; and pairwise ranking accuracy ($acc_{r}$), which measures whether the image-sentence alignment score is greater for a correct image-text pair than for its foiled pair. $acc_{r}$ is more permissive than $acc$ as it consider the model prediction correct if the score for a foil is lower than the score for a caption.
Table 22: Comparison between LLaVA-1.5-7B, TFI-7B, and FTBI-7B on the VALSE benchmark.
Existence Plurality Counting_hard Counting_small
$acc_{r}$ LLaVA-1.5-7B 69.9 13.4 35.9 35.6
TFI-7B 74.1 9.3 38.0 40.9
FTBI-7B 70.5 17.6 46.1 51.6
$acc$ LLaVA-1.5-7B 84.0 56.2 64.6 66.9
TFI-7B 85.9 54.6 66.1 68.7
FTBI-7B 84.1 58.1 71.1 74.4
$min(p_{c},p_{f})$ LLaVA-1.5-7B 73.7 16.0 52.1 45.7
TFI-7B 77.6 11.5 57.3 51.3
FTBI-7B 71.5 22.1 69.5 65.3
Counting_adversarial Relations Action Replacement Actant Swap
$acc_{r}$ LLaVA-1.5-7B 25.2 4.7 34.3 10.3
TFI-7B 24.0 2.4 29.9 11.2
FTBI-7B 36.3 8.2 37.4 19.2
$acc$ LLaVA-1.5-7B 55.6 52.0 66.4 53.1
TFI-7B 55.1 50.9 64.4 55.0
FTBI-7B 64.8 53.4 67.6 57.3
$min(p_{c},p_{f})$ LLaVA-1.5-7B 39.8 7.5 43.8 16.7
TFI-7B 41.2 4.7 35.7 16.5
FTBI-7B 59.5 14.6 52.9 30.2
Coreference Coreference_hard Foil_it
$acc_{r}$ LLaVA-1.5-7B 5.2 4.8 50.5
TFI-7B 3.1 3.9 56.8
FTBI-7B 20.2 18.3 63.0
$acc$ LLaVA-1.5-7B 52.3 52.4 75.1
TFI-7B 51.3 51.4 78.4
FTBI-7B 58.6 55.8 81.4
$min(p_{c},p_{f})$ LLaVA-1.5-7B 6.4 4.8 53.5
TFI-7B 3.8 3.9 58.3
FTBI-7B 24.0 20.2 66.6
@@
Due to the inability of LLaVA-1.5 and our models to directly output “cross_relationship_score” as the image-sentence alignment score like models such as LXMERT, we modify the computation of $acc_{r}$, $acc$, $p_{c}$ and $p_{f}$ following the approach outlined in
“lxmert_valse_eval.py” (https://github.com/Heidelberg-NLP/VALSE/
blob/main/lxmert_valse_eval.py) as follows:
@@
The results are presented in Table 22. It can be observed that TFI-7B performs better than LLaVA-1.5-7B in some areas, while FTBI-7B outperforms LLaVA-1.5-7B in all aspects, which indicates that the models infused with textual detection information are more sensitive to foiled instances and have better capabilities in visual grounding. Moreover, FTBI-7B outperforms TFI-7B on all metrics except for the “Existence” metric, further demonstrating that fine-tuning strategies are more effective than training-free approaches in helping MLLMs understand and utilize textual detection information.